# Tensor global symmetries and the Stueckelberg mechanism for tensor fields

### v1

@@ 8 Conclusions & Outlook
Several questions remain open for future investigation. The first pertains to the description of the ’t Hooft anomaly via a minimal coupling of gauge-invariant currents to background fields. A key difficulty here lies in constructing a gauge-invariant and fully conserved magnetic current. In a broader context, theories with mixed tensor symmetries often exhibit exotic dualities [40, 47, 48], and understanding the ’t Hooft anomaly in these scenarios remains an unresolved issue. A second question concerns the construction of gauge-invariant topological defects charged under mixed-symmetry tensor global symmetries. For $p$-form symmetries, one can construct Wilson loops and higher-dimensional defects by integrating gauge-invariant currents over a codimension $p+1$ surface. However, for $(p,q)$-mixed tensor symmetries, naively integrating gauge-invariant currents over a codimension $p+1$ surface does not yield a topological object. In the specific case of graviton shift symmetry, the current approach of reducing the non-gauge-invariant $(2,1)$ current to a $2$-form gauge-invariant current and then integrating it to construct topological defects is unsatisfactory. A more direct approach, building on the gauge-invariant currents developed in this work, would be of interest. In relation to this, it would be interesting to understand the bound on dimensions where spontaneous symmetry breaking is possible in these cases, essentially an analog of the Coleman-Mermin-Wagner theorem. Finally, going beyond the linear theory is another challenge. In this respect, it is worth mentioning the recent proposal of a generalized symmetry in nonlinear gravity [27] and the generalized Stueckelberg mechanism of [49] that applies to non-Abelian 2-forms.
Acknowledgements

### v2

@@ 8 Conclusions & Outlook
Several questions remain open for future investigation. The first pertains to the description of the ’t Hooft anomaly via a minimal coupling of gauge-invariant currents to background fields. A key difficulty here lies in constructing a gauge-invariant and fully conserved magnetic current. In a broader context, theories with mixed tensor symmetries often exhibit exotic dualities [49, 56, 57], and understanding the ’t Hooft anomaly in these scenarios remains an unresolved issue. A second question concerns the construction of gauge-invariant topological defects charged under mixed-symmetry tensor global symmetries. For $p$-form symmetries, one can construct Wilson loops and higher-dimensional defects by integrating gauge-invariant currents over a codimension $p+1$ surface. However, for $(p,q)$-mixed tensor symmetries, naively integrating gauge-invariant currents over a codimension $p+1$ surface does not yield a topological object. In the specific case of graviton shift symmetry, the current approach of reducing the non-gauge-invariant $(2,1)$ current to a $2$-form gauge-invariant current and then integrating it to construct topological defects is unsatisfactory. A more direct approach, building on the gauge-invariant currents developed in this work, would be of interest. In relation to this, it would be interesting to understand the bound on dimensions where spontaneous symmetry breaking is possible in these cases, essentially an analog of the Coleman-Mermin-Wagner theorem. Finally, going beyond the linear theory is another challenge. In this respect, it is worth mentioning the recent proposal of a generalized symmetry in nonlinear gravity [29, 58] and the generalized Stueckelberg mechanism of [59] that applies to non-Abelian 2-forms.
Acknowledgments